# READ-PVLA: Recurrent Adapter with Partial Video-Language Alignment for Parameter-Efficient Transfer Learning in Low-Resource Video-Language Modeling

### v1

@@ 1 $\rightarrow\frac{M}{2}$
Training parameter efficiency. The bottleneck dimension $k$ of our READ module controls how many parameters are introduced to the pre-trained backbones. Higher dimensions provide more parameters with higher risk of overfitting. In this section, we investigate such effect on the recurrent adapter and denote the results in Table 4.3. As can be seen, the performance consistently improves when the dimension increases up to 4 and saturates after this point. We note that our READ-PVLA framework can accomplish a decent performance even when the dimension reduces to one, i.e. we achieve a ROUGE-2 score of 18.57, verifying the parameter-efficiency of our proposed method.
Effects of the recurrent architecture. In addition to RNN, there exist various recurrent architectures in the literature, particularly the gated recurrent unit (GRU) (Cho et al. 2014) and long short-term memory (LSTM) (Hochreiter and Schmidhuber 1997). We experiment with different recurrent choices and explicate the results in Table 4.3. As can be observed, the performance is insensitive to the choice of recurrent design. Therefore, we select the simplest option, i.e. recurrent neural network (RNN) for our READ layers.
READ position. By default, we insert our READ layer to every Transformer block. In this ablation, we demonstrate the impact of using fewer recurrent adapters. Table 9 shows that whereas more injections of READ are more effective, the placement of READ at deeper blocks brings higher performance boost than at shallower ones. This discovery can become valuable if we have tight resource and providing a READ layer for every block is expensive. Furthermore, we denote the performance when adapting READ to various positions within a Transformer block. Table 10 reveals that after feedforward is the optimal position for our READ layer, whose observation aligns with previous works (Bapna, Arivazhagan, and Firat 2019; Pfeiffer et al. 2020).
Distance methods. We further ablate on the distance metrics to estimate the distance between video and language distributions. Technically, we perform the average- and max-pooling of the video and language representations. Then, we consider the cosine distance or the L2 distance of the two pooled vectors as the video-language distance. Results in Table 11 substantiate the superiority of our POT distance for the PVLA objective. Such success illustrates the POT-based PVLA’s advantage of modeling the relationship nature between video and language representations.
Table 8: Recurrent architecture ablation experiment on YouTube Highlights and How2. We color the settings we implement for our READ-PVLA method.
Table 7: Bottleneck dimension ablation experiments on YouTube Highlights and How2. We color the settings we implement for our READ-PVLA method.

### v2

@@ Introduction
Effects of the recurrent architecture. In addition to RNN, there exist various recurrent architectures in the literature, particularly the gated recurrent unit (GRU) (Cho et al. 2014) and long short-term memory (LSTM) (Hochreiter and Schmidhuber 1997). We experiment with different recurrent choices and explicate the results in Table 8. As can be observed, the performance is insensitive to the choice of recurrent design. Therefore, we select the simplest option, i.e. recurrent neural network (RNN) for our READ layers.
Distance methods. We further ablate on the distance metrics to estimate the distance between video and language distributions. Technically, we perform the average- and max-pooling of the video and language representations. Then, we consider the cosine distance or the L2 distance of the two pooled vectors as the video-language distance. Results in Table 9 substantiate the superiority of our POT distance for the PVLA objective. Such success illustrates the POT-based PVLA’s advantage of modeling the relationship nature between video and language representations
Qualitative Assessment
Case study. We display a TLG example on the YouTube Highlights dataset, along with the POT distance estimated by our PVLA framework and the AP score in Table 7. We observe that when the language query semantically corresponds to a moment in the video, i.e. a girl expression after she gets the proposal, the POT distance is small and correlates with the high value of AP. In contrast, when we replace the original query with an out-of-distribution one, the POT distance burgeons significantly, causing the AP to decrease from 90.28% to 30.00%. Therefore, we conclude that our READ framework is capable of intelligently adjusting the information flowing through the READ layers in order to produce the final output consistent with the video-language input and downstream tasks.
Conclusion
We propose a novel READ-PVLA framework for parameter-efficient transfer learning to video-language modeling tasks. Our READ-PVLA utilizes recurrent computation component to enable temporal modeling capability and partial video-language alignment objective to preserve critical information for bottleneck adaptation modules. Experiments demonstrate that READ-PVLA consistently outperforms both the full fine-tuning and competitive strategies, whilst bringing the benefit of parameter-efficiency (at most 1.20% trainable parameters). Our method is also applicable to diverse pre-trained models, which has the potential to employ more powerful video-language models in the future.
Acknowledgements
This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-PhD-2023-08-051T). Thong Nguyen is supported by a Google Ph.D. Fellowship in Natural Language Processing.